OpenStudy (anonymous):

Please solve 2x+4y=2800

OpenStudy (anonymous):

\[If \ 2x+4y=2800, \ then \ \\2x=2800-4y, \ and \\x=1400-2y\]

OpenStudy (anonymous):

thank you so much :)